# Double Fairness Policy Learning: Integrating Action Fairness and Outcome Fairness in Decision-making

### v1

@@ Abstract
Fairness is a central pillar of trustworthy machine learning, especially in domains where accuracy- or profit-driven optimization is insufficient. While most fairness research focuses on supervised learning, fairness in policy learning remains less explored. Because policy learning is interventional, it induces two distinct fairness targets: action fairness (equitable action assignments) and outcome fairness (equitable downstream consequences). Crucially, equalizing actions does not generally equalize outcomes when groups face different constraints or respond differently to the same action. We propose a novel double fairness learning (DFL) framework that explicitly manages the trade-off among three objectives: action fairness, outcome fairness, and value maximization. We integrate fairness directly into a multi-objective optimization problem for policy learning and employ a lexicographic weighted Tchebyshev method that recovers Pareto solutions beyond convex settings, with theoretical guarantees on the regret bounds. Our framework is flexible and accommodates various commonly used fairness notions. Extensive simulations demonstrate improved performance relative to competing methods. In applications to a motor third-party liability insurance dataset and an entrepreneurship training dataset, DFL substantially improves both action and outcome fairness while incurring only a modest reduction in overall value.
Keywords: Fairness; Trustworthiness; Policy Learning; Multi-objective Optimization.
@@ 1 Introduction
As a core aspect of trustworthiness, fairness has become central to machine learning (Dwork et al., 2012; Zemel et al., 2013; Grgic-Hlaca et al., 2016; Hardt et al., 2016; Kusner et al., 2017; Wu et al., 2019; Tan et al., 2022; Chen et al., 2023), motivated by ethical concerns about the societal impacts of algorithmic decision-making. In many domains such as healthcare, insurance, lending, and public programs, algorithmic systems can affect large populations at scale. Purely accuracy- or profit-driven optimization can amplify disparities, imposing disproportionate burdens on vulnerable groups, and exposing organizations to legal and reputational risk. Fairness is therefore not merely a normative ideal. It is an operational requirement for responsible deployment and a critical pillar of ethical AI.
Most existing work on algorithmic fairness has focused on supervised learning. In contrast, fairness in policy learning, where the goal is to maximize expected reward or social welfare, has received comparatively limited attention (Nabi et al., 2019; Tan et al., 2022; Viviano and Bradic, 2023; Fang et al., 2023; Wang et al., 2025). A key distinction is that policy learning is inherently interventional: a policy selects an action $A$, and that action alters subsequent outcomes. This leads to two conceptually distinct fairness targets. Action fairness concerns whether the policy treats individuals similarly in its action choices across sensitive groups (e.g., equalized treatment assignments). Outcome fairness instead asks whether the consequences (e.g., the benefits, welfare, or other fairness-relevant outcomes) induced by the policy are equitably distributed across groups. Importantly, equalizing actions does not, by itself, equalize outcomes when groups face different constraints, environments, or downstream responses to the same action. Consistent with this distinction, our analysis of a Belgian motor liability insurance dataset (Dutang and Charpentier, 2019) show that standard policy optimization methods can induce substantial disparities in both action and outcome fairness (Table 1 in Section 6).
@@
This work is also motivated by growing concerns about fairness in the insurance industry. Machine learning algorithms are increasingly used to automate risk assessment, pricing, and claims-related decisions, making trustworthy decision rules especially critical. For example, UnitedHealth has used the nH Predict platform (Talia, 2024) to assess patient risk and forecast outcomes in ways that inform coverage and pricing. While such tools can improve efficiency, they also raise concerns about algorithmic bias and discrimination against vulnerable populations. A recent class-action lawsuit alleges that the nH Predict algorithm was used to unlawfully deny or curtail rehabilitation care for seriously ill and elderly. [^1] See the news report at Healthcare Finance News.
@@
The DFL framework raises three central challenges. First, outcome fairness can only be controlled indirectly through the policy’s actions and the environment’s response: unlike action fairness, one cannot directly “intervene” on outcomes, so it is unclear when an outcome-fair (or doubly fair) policy exists at all. Second, the three objectives, action fairness, outcome fairness and value maximization, typically conflict, so it is rare for a single policy to optimize all three simultaneously. Hence, a principled framework is needed to characterize and manage these trade-offs, for which we adopt the concept of Pareto optimality. Third, computing the relevant Pareto fairness set is nontrivial in practice: standard linear scalarization approaches can miss Pareto-optimal solutions when the policy class is non-convex, as is typical for many policy parameterizations (Schulman et al., 2015).
We address these challenges with a unified methodology and theory. We first provide a systematic analysis of when outcome-fair and doubly fair policies can be achieved, yielding interpretable conditions for feasibility. To manage trade-offs, we cast DFL as a multi-objective optimization problem (Miettinen, 1999) and target policies that are Pareto-efficient with respect to the two fairness criteria while achieving high value function. To compute Pareto solutions, we adopt a lexicographic weighted Tchebyshev scalarization strategy (Steuer and Choo, 1983), which can recover Pareto solutions without requiring convexity (see Figure 7 in Appendix). We then develop an estimator that embeds fairness objectives directly into the policy optimization procedure, establish regret bounds that provide rigorous performance guarantees under double fairness constraints.
@@
Related Literature. Policy learning seeks to maximize expected reward by learning an optimal policy that selects the best action for each situation. It has broad applications in precision medicine (Murphy, 2003; Robins, 2004; Qian and Murphy, 2011; Zhao et al., 2012; Wallace and Moodie, 2015; Wang et al., 2018; Shi et al., 2018; Qi et al., 2020; Cui and Tchetgen Tchetgen, 2021; Bian et al., 2023), economics (Manski, 2004; Kitagawa and Tetenov, 2018; Athey and Wager, 2021), and business strategy (Den Boer, 2015; Javanmard and Nazerzadeh, 2019; Bian et al., 2024). This literature largely focuses on value maximization, whereas our work augments policy learning with two complementary fairness targets—action fairness (fairness of decisions) and outcome fairness (fairness of the induced consequences)—and studies how to optimize value under both.
Work on fairness-aware policy learning is comparatively scarce and typically enforces a single fairness notion. Nabi et al. (2019) formulate fair policy learning via causal mediation analysis, specifying impermissible causal pathways from the sensitive attribute and optimizing policies under the resulting causal constraints. Tan et al. (2022), Fang et al. (2023), and Cui and Han (2025) develop quantile-optimal decision rules designed to protect disadvantaged groups by controlling tail outcomes. Viviano and Bradic (2023) adopt a “first do no harm” principle by characterizing a welfare Pareto frontier and selecting the fairest allocation within that welfare-efficient set, with computation based on mixed-integer optimization. Wang et al. (2025) study counterfactual fairness in sequential decision making, proposing approaches tailored to reinforcement learning settings. In contrast to these works, we explicitly distinguish action fairness from outcome fairness and address the double fairness problem with different theoretical tools. Figure 1 compares our method to existing fairness-aware policy learning approaches. We group prior single-fairness methods into two categories: action single fair (ASF) methods, which enforce only action fairness, and outcome single fair (OSF) methods, which enforce only outcome fairness. Figure 1 shows that our proposed DFL framework achieves consistently strong performance across all three objectives (action fairness, outcome fairness, and value) with modest compromises in each. By contrast, competing approaches typically improve one dimension at the expense of substantial degradation in the others.
@@ 2 Preliminary
Throughout, we allow for a general two-outcome formulation : $R\equiv(R^{(1)},R^{(2)})^{\top}\in\mathbb{R}^{2}$, where $R^{(1)}$ is the primary outcome that the decision maker seeks to maximize, while $R^{(2)}$ is an outcome along which fairness or equity considerations are assessed. This formulation accommodates a range of applications. In the CASdatasets (Section 6.1), $R^{(1)}$ represents company revenue, while $R^{(2)}$ denotes customer welfare. We also accommodate the degenerate case where $R^{(1)}=R^{(2)}$: for instance, in the government job-training data (Section 6.2), both components correspond to subsequent entrepreneurial activity. In dynamic pricing, $R^{(1)}$ typically represents net profit, whereas $R^{(2)}$ might denote demand or consumer surplus—metrics for which fairness has been extensively studied (see, e.g., Cohen et al., 2022; Bian et al., 2026). Furthermore, in healthcare, $R^{(1)}$ may represent clinical outcomes, while $R^{(2)}$ denotes the allocation of medical resources.
@@ 2.1 Policy Learning
where $\widehat{r}(s,x,a)$ is the regression function estimator of the reward function $r(s,x,a)$. Such a regression approach has been widely used in the policy learning literature; see, e.g., Qian and Murphy (2011); Moodie et al. (2012). Then in the value maximization step, the estimated optimal policy is derived by maximizing this estimated value function, within the pre-specified policy class $\Pi$:
@@ 2.2 Fairness Notions
Recently, many fairness notions have been proposed under supervised learning setting, with particular attention to equal opportunity (Hardt et al., 2016) and counterfactual fairness (Kusner et al., 2017). Here, we extend these two notions to the policy learning framework and formally define both action fairness and outcome fairness under each concept.
@@
Equal opportunity requires that individuals with the same observed information $X$ be treated equally across different sensitive groups. This criterion is intuitively appealing and has been widely used in practice (Hardt et al., 2016). However, because it conditions only on observed covariates, it may fail to account for historical or unobserved sources of bias (Chen et al., 2023). To address this, Kusner et al. (2017) proposed counterfactual fairness, which adopts a causal perspective by asking whether the decision (or the induced outcome) would have remained unchanged had the individual belonged to a different sensitive group.
@@ 3.2 Multi-objective Policy Learning
Making optimal decisions in the presence of trade-offs among multiple objectives is inherently challenging, as it is generally impossible to optimize all objectives simultaneously. Therefore, it is crucial to understand how to effectively balance multiple objectives and make optimal trade-offs among them. The concept of Pareto set (see, e.g., Miettinen, 1999) offers a valuable framework for this purpose, capturing the idea that improvements in one objective cannot be achieved without compromising at least one other. By leveraging the idea of Pareto set, we can navigate complex decision-making scenarios and strive for policies that optimize overall benefits across diverse objectives.
@@
Proposition 3 (Lexicographic Weighted Tchebychef Method (Steuer and Choo, 1983)).
@@
Viviano and Bradic (2023) adopted the linear scalarization approach (specific formulation discussed in Section 5) to characterize the Pareto set. However, this method cannot identify Pareto fairness policies in the non-convex regions of the Pareto fairness front (as illustrated in Figure 7 in Section 8.1 of the Supplemental Materials). By contrast, the lexicographic weighted Tchebycheff approach can identify the full set of Pareto fairness policies without requiring convexity. We also include a simple intuitive example in Section 8.1 to demonstrate that the linear scalarization approach fails, while the weighted Tchebycheff method works.
@@
7:  Compute $k^{*}=\operatorname*{arg\,max}_{1\leq k\leq K}\widehat{V}_{k}^{*}$ .
8:  return $\widehat{\pi}=\widehat{\pi}_{k^{*}}^{*}$.
@@ 4 Theoretical Results
Theorem 1 shows the regret consists of two parts: (i) the uniform policy evaluation error over $\Pi_{p}$ and $\Pi$, and (ii) the estimation error of the Pareto fairness set, measured by the Hausdorff distance $H_{\mathrm{TV}}(\Pi_{p},\widehat{\Pi}_{p})$. For (i), if the nuisance functions are estimated with sufficient accuracy and $\Pi$ is a VC subgraph class, then the uniform error admits a suitable bound; see, e.g., Kitagawa and Tetenov (2018); Athey and Wager (2021). We now analyze component (ii), starting with the following assumption.
@@
Assumption 4.2 imposes a separation condition: policies that are at least $\delta$ away from $\Pi_{p}$ in total variation must incur an objective gap of order $\delta$ under either the criterion $\Delta(\cdot)$ or the weighted Tchebycheff criterion $M_{\alpha}(\cdot)$. Such well-separated solution conditions are standard in establishing consistency of $M$- and $Z$-estimators; see, e.g., Van der Vaart (2000).
@@ 4.1 Equal Opportunity and Counterfactual Fairness Applications
Assumption 4.3 is commonly imposed in policy learning literature, see, e.g., Kitagawa and Tetenov (2018); Athey and Wager (2021). Common examples of classes with finite VC-subgraph dimension include linear decision rules, decision trees, and sigmoid (logistic) score classes.
@@
Assumption 4.4 is needed because the outcome fairness metric estimator $\widehat{\Delta}_{2}(\pi)$ and the value estimator $\widehat{V}(\pi)$ are constructed from $\widehat{r}$ and $\widehat{r}$, respectively. The exponents $\gamma_{1}$ and $\gamma_{2}$ depend on the methods used to estimate $r$ and $f$. If $r$ and $f$ are estimated using correctly specified parametric models, one typically has $\gamma_{1}=\gamma_{2}=1$. Alternatively, flexible nonparametric methods such as sieve estimators (Huang, 1998) can be used. Under standard smoothness and complexity conditions, sieve estimators can achieve the optimal mean-squared rate $O_{p}(n^{-\frac{2\beta}{2\beta+d}})$, where $\beta$ is the smoothness parameter of the reward function, and $d$ is the dimension of $X$. Sieve methods are widely used in the policy evaluation and learning literature; see, e.g., Shi et al. (2022); Chen and Qi (2022); Bian et al. (2025).
For the counterfactual fairness notion, we additionally impose the following Assumption 4.5 to ensure that the counterfactual covariates $X_{s,x}(s^{\prime})$ are identifiable from the observed data. Similar assumptions are used in Chen et al. (2023) and Wang et al. (2025) to study counterfactual fairness.
@@
Assumption 4.6(i) is standard and has also been imposed by Wang et al. (2025) for counterfactual fairness policy learning. Assumption 4.6(ii) requires only Lipschitz continuity in expected mean square, which is mild enough to accommodate deterministic policies (for which pointwise Lipschitz continuity typically fails). A closely related condition is used by Shi et al. (2021) when analyzing linear deterministic decision rules; see Assumption A5(ii) therein.
@@
$|\mu(\widehat{\pi})|=O_{p}\big(\sqrt{\mbox{VC}(\Pi)/n}+\sqrt{\log n/n}+\sqrt{d/n}\big)++O_{p}\big(n^{-\frac{\gamma_{1}}{2}}+n^{-\frac{\gamma_{2}}{2}}\big).$
@@
In Theorems 3 and 4, the slack parameter $\kappa$ involves the term $n^{-\frac{\gamma_{2}}{2}}$. In practice, if a parametric model is used, one can typically choose $\gamma_{2}=1$. For nonparametric regression methods such as the sieve estimator, $\gamma_{2}$ is given by $\frac{2\beta}{2\beta+d}$, as noted earlier. Since $\beta$ is typically unknown, one may infer it from the optimal complexity choice. For sieve estimation, the optimal number of basis functions satisfies $L^{\mathrm{opt}}\;=\;n^{\frac{d}{d+2\beta}}.$ Thus, one can first select $L^{\text{opt}}$ (e.g., via cross-validation), and then back out $\beta$ and hence $\gamma_{2}$ accordingly. The same logic extends to other nonparametric methods (e.g., kernel or local polynomial estimators), whose optimal bandwidth choices also depend on $\beta$; see, for example, Tsybakov (2009).

### v2

@@ Abstract
Fairness has recently received significant attention in machine learning, driven by ethical concerns surrounding algorithmic applications. However, fairness in policy learning, where the goal is to make fair decision while simultaneously maximizing expected rewards, has been less explored. Most existing works in fairness for policy learning focus primarily on action fairness. However, many applications require fairness in both the actions taken and their resulting outcomes. This paper tackles the challenge of double fairness in policy learning by considering the trade-off of three objectives simultaneously: action fairness, outcome fairness, and maximizing the value function. We propose a novel framework that integrates fairness directly into a multi-objective optimization problem for policy learning. Our framework is flexible and accommodates various commonly used fairness notions. We establish theoretical guarantees for the proposed method and demonstrate its empirical superiority over competing approaches. Finally, we apply our method to a motor third-party liability insurance dataset and an entrepreneurship training dataset. Our proposed method demonstrates greater fairness on both fairness metrics while having only a modest reduction in overall revenue.
keywords:
Fairness; Trustworthiness; Policy Learning; Multi-objective Optimization
@@ 1 Introduction
As a core aspect of trustworthiness, fairness has become central to machine learning (dwork2012fairness; zemel2013learning; grgic2016case; hardt2016equality; kusner2017counterfactual; wu2019counterfactual; tan2022rise; chen2023learning), motivated by ethical concerns about the societal impacts of algorithmic decision-making. In many domains such as healthcare, insurance, lending, and public programs, algorithmic systems can affect large populations at scale. Purely accuracy- or profit-driven optimization can amplify disparities, imposing disproportionate burdens on vulnerable groups, and exposing organizations to legal and reputational risk. Fairness is therefore not merely a normative ideal. It is an operational requirement for responsible deployment and a critical pillar of ethical AI.
Most existing work on algorithmic fairness has focused on supervised learning. In contrast, fairness in policy learning, where the goal is to maximize expected reward or social welfare, has received comparatively limited attention (nabi2019learning; tan2022rise; viviano2023fair; fang2023fairness; wang2025counterfactually). A key distinction is that policy learning is inherently interventional: a policy selects an action $A$, and that action alters subsequent outcomes. This leads to two conceptually distinct fairness targets. Action fairness concerns whether the policy treats individuals similarly in its action choices across sensitive groups (e.g., equalized treatment assignments). Outcome fairness instead asks whether the consequences (e.g., the benefits, welfare, or other fairness-relevant outcomes) induced by the policy are equitably distributed across groups. Importantly, equalizing actions does not, by itself, equalize outcomes when groups face different constraints, environments, or downstream responses to the same action. Consistent with this distinction, our analysis of a Belgian motor liability insurance dataset (dutang2019casdatasets) show that standard policy optimization methods can induce substantial disparities in both action and outcome fairness (Table 1 in Section 6).
@@
This work is also motivated by growing concerns about fairness in the insurance industry. Machine learning algorithms are increasingly used to automate risk assessment, pricing, and claims-related decisions, making trustworthy decision rules especially critical. For example, UnitedHealth has used the nH Predict platform (talia2024algorithms) to assess patient risk and forecast outcomes in ways that inform coverage and pricing. While such tools can improve efficiency, they also raise concerns about algorithmic bias and discrimination against vulnerable populations. A recent class-action lawsuit alleges that the nH Predict algorithm was used to unlawfully deny or curtail rehabilitation care for seriously ill and elderly. [^1] See the news report at Healthcare Finance News.
@@
The DFL framework raises three central challenges. First, outcome fairness can only be controlled indirectly through the policy’s actions and the environment’s response: unlike action fairness, one cannot directly “intervene” on outcomes, so it is unclear when an outcome-fair (or doubly fair) policy exists at all. Second, the three objectives, action fairness, outcome fairness and value maximization, typically conflict, so it is rare for a single policy to optimize all three simultaneously. Hence, a principled framework is needed to characterize and manage these trade-offs, for which we adopt the concept of Pareto optimality. Third, computing the relevant Pareto fairness set is nontrivial in practice: standard linear scalarization approaches can miss Pareto-optimal solutions when the policy class is non-convex, as is typical for many policy parameterizations (schulman2015trust).
We address these challenges with a unified methodology and theory. We first provide a systematic analysis of when outcome-fair and doubly fair policies can be achieved, yielding interpretable conditions for feasibility. To manage trade-offs, we cast DFL as a multi-objective optimization problem (miettinen1999nonlinear) and target policies that are Pareto-efficient with respect to the two fairness criteria while achieving high value function. To compute Pareto solutions, we adopt a lexicographic weighted Tchebyshev scalarization strategy (steuer1983interactive), which can recover Pareto solutions without requiring convexity (see Figure 7 in Appendix). We then develop an estimator that embeds fairness objectives directly into the policy optimization procedure, establish regret bounds that provide rigorous performance guarantees under double fairness constraints.
@@
Related Literature. Policy learning seeks to maximize expected reward by learning an optimal policy that selects the best action for each situation. It has broad applications in precision medicine (Murphy; gest; qian2011performance; zhao2012estimating; wallace2015doubly; wang2018quantile; shi2018high; qi2020multi; cui2021semiparametric; bian2023variable), economics (manski2004statistical; kitagawa2018should; athey2021policy), and business strategy (den2015dynamic; javanmard2019dynamic; bian2024tale). This literature largely focuses on value maximization, whereas our work augments policy learning with two complementary fairness targets—action fairness (fairness of decisions) and outcome fairness (fairness of the induced consequences)—and studies how to optimize value under both.
Work on fairness-aware policy learning is comparatively scarce and typically enforces a single fairness notion. nabi2019learning formulate fair policy learning via causal mediation analysis, specifying impermissible causal pathways from the sensitive attribute and optimizing policies under the resulting causal constraints. tan2022rise, fang2023fairness, and cui2025policy develop quantile-optimal decision rules designed to protect disadvantaged groups by controlling tail outcomes. viviano2023fair adopt a “first do no harm” principle by characterizing a welfare Pareto frontier and selecting the fairest allocation within that welfare-efficient set, with computation based on mixed-integer optimization. wang2025counterfactually study counterfactual fairness in sequential decision making, proposing approaches tailored to reinforcement learning settings. In contrast to these works, we explicitly distinguish action fairness from outcome fairness and address the double fairness problem with different theoretical tools. Figure 1 compares our method to existing fairness-aware policy learning approaches. We group prior single-fairness methods into two categories: action single fair (ASF) methods, which enforce only action fairness, and outcome single fair (OSF) methods, which enforce only outcome fairness. Figure 1 shows that our proposed DFL framework achieves consistently strong performance across all three objectives (action fairness, outcome fairness, and value) with modest compromises in each. By contrast, competing approaches typically improve one dimension at the expense of substantial degradation in the others.
@@ 2 Preliminary
Throughout, we allow for a general two-outcome formulation : $R\equiv(R^{(1)},R^{(2)})^{\top}\in\mathbb{R}^{2}$, where $R^{(1)}$ is the primary outcome that the decision maker seeks to maximize, while $R^{(2)}$ is an outcome along which fairness or equity considerations are assessed. This formulation accommodates a range of applications. In the CASdatasets (Section 6.1), $R^{(1)}$ represents company revenue, while $R^{(2)}$ denotes customer welfare. We also accommodate the degenerate case where $R^{(1)}=R^{(2)}$: for instance, in the government job-training data (Section 6.2), both components correspond to subsequent entrepreneurial activity. In dynamic pricing, $R^{(1)}$ typically represents net profit, whereas $R^{(2)}$ might denote demand or consumer surplus—metrics for which fairness has been extensively studied (see, e.g., cohen2022price; bian2026beyond). Furthermore, in healthcare, $R^{(1)}$ may represent clinical outcomes, while $R^{(2)}$ denotes the allocation of medical resources.
@@ 2.1 Policy Learning
where $\widehat{r}(s,x,a)$ is the regression function estimator of the reward function $r(s,x,a)$. Such a regression approach has been widely used in the policy learning literature; see, e.g., qian2011performance; moodie2012q. Then in the value maximization step, the estimated optimal policy is derived by maximizing this estimated value function, within the pre-specified policy class $\Pi$:
@@ 2.2 Fairness Notions
Recently, many fairness notions have been proposed under supervised learning setting, with particular attention to equal opportunity (hardt2016equality) and counterfactual fairness (kusner2017counterfactual). Here, we extend these two notions to the policy learning framework and formally define both action fairness and outcome fairness under each concept.
@@
Equal opportunity requires that individuals with the same observed information $X$ be treated equally across different sensitive groups. This criterion is intuitively appealing and has been widely used in practice (hardt2016equality). However, because it conditions only on observed covariates, it may fail to account for historical or unobserved sources of bias (chen2023learning). To address this, kusner2017counterfactual proposed counterfactual fairness, which adopts a causal perspective by asking whether the decision (or the induced outcome) would have remained unchanged had the individual belonged to a different sensitive group.
@@ 3.2 Multi-objective Policy Learning
Making optimal decisions in the presence of trade-offs among multiple objectives is inherently challenging, as it is generally impossible to optimize all objectives simultaneously. Therefore, it is crucial to understand how to effectively balance multiple objectives and make optimal trade-offs among them. The concept of Pareto set (see, e.g., miettinen1999nonlinear) offers a valuable framework for this purpose, capturing the idea that improvements in one objective cannot be achieved without compromising at least one other. By leveraging the idea of Pareto set, we can navigate complex decision-making scenarios and strive for policies that optimize overall benefits across diverse objectives.
@@
Proposition 3 (Lexicographic Weighted Tchebychef Method (steuer1983interactive)).
@@
viviano2023fair adopted the linear scalarization approach (specific formulation discussed in Section 5) to characterize the Pareto set. However, this method cannot identify Pareto fairness policies in the non-convex regions of the Pareto fairness front (as illustrated in Figure 7 in Section 8.1 of the Supplemental Materials). By contrast, the lexicographic weighted Tchebycheff approach can identify the full set of Pareto fairness policies without requiring convexity. We also include a simple intuitive example in Section 8.1 to demonstrate that the linear scalarization approach fails, while the weighted Tchebycheff method works.
@@
7:  end for
8:  Compute $k^{*}=\operatorname*{arg\,max}_{1\leq k\leq K}\widehat{V}_{k}^{*}$ .
9:  return $\widehat{\pi}=\widehat{\pi}_{k^{*}}^{*}$.
10:end function
@@ 4 Theoretical Results
Theorem 1 shows the regret consists of two parts: (i) the uniform policy evaluation error over $\Pi_{p}$ and $\Pi$, and (ii) the estimation error of the Pareto fairness set, measured by the Hausdorff distance $H_{\mathrm{TV}}(\Pi_{p},\widehat{\Pi}_{p})$. For (i), if the nuisance functions are estimated with sufficient accuracy and $\Pi$ is a VC subgraph class, then the uniform error admits a suitable bound; see, e.g., kitagawa2018should; athey2021policy. We now analyze component (ii), starting with the following assumption.
@@
Assumption 4.2 imposes a separation condition: policies that are at least $\delta$ away from $\Pi_{p}$ in total variation must incur an objective gap of order $\delta$ under either the criterion $\Delta(\cdot)$ or the weighted Tchebycheff criterion $M_{\alpha}(\cdot)$. Such well-separated solution conditions are standard in establishing consistency of $M$- and $Z$-estimators; see, e.g., van2000asymptotic.
@@ 4.1 Equal Opportunity and Counterfactual Fairness Applications
Assumption 4.3 is commonly imposed in policy learning literature, see, e.g., kitagawa2018should; athey2021policy. Common examples of classes with finite VC-subgraph dimension include linear decision rules, decision trees, and sigmoid (logistic) score classes.
@@
Assumption 4.4 is needed because the outcome fairness metric estimator $\widehat{\Delta}_{2}(\pi)$ and the value estimator $\widehat{V}(\pi)$ are constructed from $\widehat{r}$ and $\widehat{r}$, respectively. The exponents $\gamma_{1}$ and $\gamma_{2}$ depend on the methods used to estimate $r$ and $f$. If $r$ and $f$ are estimated using correctly specified parametric models, one typically has $\gamma_{1}=\gamma_{2}=1$. Alternatively, flexible nonparametric methods such as sieve estimators (huang1998projection) can be used. Under standard smoothness and complexity conditions, sieve estimators can achieve the optimal mean-squared rate $O_{p}(n^{-\frac{2\beta}{2\beta+d}})$, where $\beta$ is the smoothness parameter of the reward function, and $d$ is the dimension of $X$. Sieve methods are widely used in the policy evaluation and learning literature; see, e.g., shi2020statistical; chen2022well; bian2025off.
For the counterfactual fairness notion, we additionally impose the following Assumption 4.5 to ensure that the counterfactual covariates $X_{s,x}(s^{\prime})$ are identifiable from the observed data. Similar assumptions are used in chen2023learning and wang2025counterfactually to study counterfactual fairness.
@@
Assumption 4.6(i) is standard and has also been imposed by wang2025counterfactually for counterfactual fairness policy learning. Assumption 4.6(ii) requires only Lipschitz continuity in expected mean square, which is mild enough to accommodate deterministic policies (for which pointwise Lipschitz continuity typically fails). A closely related condition is used by shi2021concordance when analyzing linear deterministic decision rules; see Assumption A5(ii) therein.
@@
$|\mu(\widehat{\pi})|=O_{p}\big(\sqrt{\mbox{VC}(\Pi)/n}+\sqrt{\log n/n}+\sqrt{d/n}\big)+O_{p}\big(n^{-\frac{\gamma_{1}}{2}}+n^{-\frac{\gamma_{2}}{2}}\big).$
@@
In Theorems 3 and 4, the slack parameter $\kappa$ involves the term $n^{-\frac{\gamma_{2}}{2}}$. In practice, if a parametric model is used, one can typically choose $\gamma_{2}=1$. For nonparametric regression methods such as the sieve estimator, $\gamma_{2}$ is given by $\frac{2\beta}{2\beta+d}$, as noted earlier. Since $\beta$ is typically unknown, one may infer it from the optimal complexity choice. For sieve estimation, the optimal number of basis functions satisfies $L^{\mathrm{opt}}\;=\;n^{\frac{d}{d+2\beta}}.$ Thus, one can first select $L^{\text{opt}}$ (e.g., via cross-validation), and then back out $\beta$ and hence $\gamma_{2}$ accordingly. The same logic extends to other nonparametric methods (e.g., kernel or local polynomial estimators), whose optimal bandwidth choices also depend on $\beta$; see, for example, Tsybakov2009.